### v3

@@ 7.2 MGS for Quasigroups
    The decision version belongs to $\exists^{\log^{2}n}\forall^{\log n}\mathsf{NTIME}(\mathrm{polylog}(n))$;
@@
    The decision version belongs to $\exists^{\log^{2}n}\mathsf{SAC}^{1}\subseteq\mathsf{DSPACE}(\log^{2}n)$; [^10] Wolf [Wol94] showed the containment $\exists^{\log^{2}n}\mathsf{SAC}^{1}\subseteq\textsf{DSPACE}(\log^{2}n)$. and
  3. c.
    The search version belongs to $\mathsf{quasiAC^{0}}$.
@@
    not $\exists^{\log^{2}n}\mathsf{P}$-complete under $\mathsf{quasiAC^{0}}$ Turing reductions.
@@
(a) Thm. 7.5 (a) for MGS, resp. Thm. 4.1 for Quasigroup Isomorphism, place both problems into classes that are contained in $\mathsf{quasiAC^{0}}$ by Fact 2.4. The result then follows by noting that $\mathsf{quasiAC^{0}}$ is closed under $\mathsf{quasiAC^{0}}$ Turing reductions, $\textsc{Parity}\in\mathsf{P}\subseteq\exists^{\log^{2}n}\mathsf{P}$, but $\textsc{Parity}\notin\mathsf{quasiAC^{0}}$.
(b) Both MGS for quasigroups and Quasigroup Isomorphism are in $\mathsf{DSPACE}(\log^{2}n)$ by Thm. 7.5 (b), resp. [CTW13]. The closure of $\mathsf{DSPACE}(\log^{2}n)$ under poly-log space reductions is contained in $\mathsf{polyL}=\bigcup_{k\geq 0}\mathsf{DSPACE}(\log^{k}n)$. If either of these two quasigroup problems were complete for $\exists^{\log^{2}n}\mathsf{P}$ under polylog-space Turing reductions, we would get $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{polyL}$. Under the latter assumption, by a straightforward padding argument, we now show that $\mathsf{EXP}=\mathsf{PSPACE}$.
@@
Now we return to establishing the main result of this section, Thm. 7.5. To establish Thm. 7.5 (a) and (c), we will crucially leverage the Membership for quasigroups problem. To this end, we will first establish the following.
@@
Membership for quasigroups belongs to $\mathsf{NTIME}(\mathrm{polylog}(n))$.
Remark 7.8.
A natural approach to Thm. 7.5 (b) is also to attempt to use Thm. 7.7. For comparison to the approach we take instead, we note that a careful analysis of the proof for Thm. 7.7 only yields a $\mathsf{DSPACE}(\log^{3}n)$ bound for MGS for quasigroups.
@@
The proofs of Theorem 7.7 and Theorem 7.5 rely crucially on the following adaption of the Babai–Szemerédi Reachability Lemma [BS84, Theorem 3.1] to quasigroups. We first generalize the notion of a straight-line program for groups [BS84] to SLPs for quasigroups. Let $X$ be a set of generators for a quasigroup $G$. We call a sequence of elements $g_{1},\ldots,g_{\ell}\in G$ a straight-line program (SLP for short) if each $g_{i}$ ($i\in[\ell]$) either belongs to $X$, or is of the form or $g_{j}g_{k}$, $g_{j}\backslash g_{k}$, or $g_{j}/g_{k}$ for some $j,k<i$ (where $g_{j}\backslash g_{k}$, resp. $g_{j}/g_{k}$, denotes the quasigroup division as defined in Section 2.1). An SLP is said to compute or generate a set $S$ (or an element $g$) if $S\subseteq\{g_{1},\dots,g_{\ell}\}$ (resp. $g\in\{g_{1},\dots,g_{\ell}\}$).
@@
Let $G$ be a finite quasigroup and let $X$ be a set of generators for $G$. For each $g\in G$, there exists a straight-line program over $X$ generating $g$ which has length $O(\log^{2}|G|)$.
@@
We follow the same strategy as in the proof of [BS84, Theorem 3.1], but there are some subtle, yet crucial, modifications due to the fact that quasigroups are non-associative and need not posses an identity element. We will inductively construct a sequence $z_{0},z_{1},\ldots,z_{t}$ (similar to a cube generating sequence). For any sequence of elements $z_{1},\dotsc,z_{k}$, let $P(z_{1}z_{2}\dotsb z_{k})$ denote the left-to-right parenthesization, e.g., $P(z_{1}z_{2}z_{3})=(z_{1}z_{2})z_{3}$. For some initial segment $z_{0},z_{1},\ldots,z_{i}$ define the cube
$K(i)=\{P(z_{0}z_{1}^{e_{1}}\cdots z_{i}^{e_{i}}):e_{1},\ldots,e_{i}\in\{0,1\}\},$
where $e_{j}=0$ denotes omitting $z_{j}$ from the product (since there need not be an identity element).
Define $L(i)=K(i)\backslash K(i)=\{g\backslash h:g,h\in K(i)\}$. We will construct a sequence $z_{0},z_{1},\ldots,z_{t}$ such that $t\leq\lceil\log_{2}(n)\rceil$ and $L(t)=G$. Moreover, we derive a bound on the straight-line cost $c(i)$ for $\{z_{0},z_{1},\ldots,z_{i}\}$ ($1\leq i\leq t$), which is defined as the length of the shortest SLP generating $\{z_{0},z_{1},\ldots,z_{i}\}$.
We take $z_{0}$ as an arbitrary element from $X$. Hence, $K(0)=\{z_{0}\}$, and so $c(0)=1$. Next, let us construct $K(i+1)$ from $K(i)$. If $L(i)\neq G$, we set $z_{i+1}$ to be the element $z^{\prime}\notin L(i)$ that minimizes $c(i+1)-c(i)$. We first claim that $|K(i+1)|=2\cdot|K(i)|$. Note that $K(i+1)=K(i)\cup K(i)z_{i+1}$ by definition. As right-multiplication by a fixed element is a bijection in a quasi-group, it suffices to show that $K(i)\cap K(i)z_{i+1}=\emptyset$. So, suppose that there exists some $a\in K(i)\cap K(i)z_{i+1}$. Then $a=gz_{i+1}$ for some $g\in K(i)$. Hence, $z_{i+1}=g\backslash a$, contradicting $z_{i+1}\not\in L(i)$, since both $a$ and $g$ are in $K(i)$.
It now follows that $\left|\mathinner{K(i)}\right|=2^{i}$ and, hence, $t\leq\lceil\log_{2}(|G|)\rceil$ and $L(t)=G$. Moreover, for every $g\in G$ we obtain an SLP of length at most $c(t)+2t+1$: write $g=a\backslash b$ for $a,b\in K(t)$ and start with the SLP computing $\{z_{0},z_{1},\ldots,z_{t}\}$. We obtain an SLP for $a=z_{0}z_{1}^{e_{1}}\cdots z_{t}^{e_{t}}$ by adding a new element for each $z_{0}z_{1}^{e_{1}}\cdots z_{i}^{e_{i}}$ with $e_{i}=1$ ($1\leq i\leq t$, thus, at most $t$ new elements). Likewise, we get an SLP for $b$. Adding a last element $g=a\backslash b$, we obtain an SLP computing $g$.
It remains to bound the straight-line cost $c(i)$ of $\{z_{0},z_{1},\ldots,z_{i}\}$. Here, we claim that that $c(i+1)-c(i)\leq 4i+3$ (note that this is slightly worse than the bound $c(i+1)-c(i)\leq 2i+1$ for groups [BS84]). It follows from this claim that $c(i)\in O(i^{2})$. We will now turn to proving our claim.
Proof of Claim.
If $G\neq L(i)$, then either $X\not\subseteq L(i)$ or $L(i)$ is not a sub-quasigroup. Hence, we have one of the following cases:
@@
The result now follows. ∎
For proving Theorem 7.7 we follow essentially the ideas of [Fle22] (though we avoid introducing the notion of Cayley circuits). Fleischer obtained a $\mathsf{quasiAC^{0}}$ bound for Membership for groups by then showing that the Cayley circuits for this problem can be simulated by a $\mathsf{quasiAC^{0}}$ circuit. We will instead directly analyze the straight-line programs using an $\mathsf{NTIME}(\text{polylog}(n))$ (sequential) algorithm.
@@
To decide whether $g\in\langle X\rangle$, we guess a sequence of elements $g_{1},\dots,g_{\ell}\in G$ with $\ell\in O(\log^{2}|G|)$ and $g_{\ell}=g$ and in a next step verify whether this, indeed, is a straight-line program. Clearly, the conditions ($g_{i}$ ($i\in[\ell]$) either belongs to $X$, or is of the form $g_{j}g_{k}$, $g_{j}\backslash g_{k}$, or $g_{j}/g_{k}$ for some $j,k<i$) can be checked in time polynomial in the length of the straight-line program and, thus, in time polylogarithmic in the input length. ∎
The proof of Thm. 7.5 (a) and (c) below is by describing a reduction from MGS to Membership. To establish Thm. 7.5 (b), we use the following result of Wagner:
Lemma 7.11 ([Wag10, Theorem 10.2.1]).
Let $Q$ be a quasigroup, and let $X\subseteq Q$. We can compute $\langle X\rangle$ in $\mathsf{SAC}^{1}$.
We now have all the ingredients we need to prove all three parts of Thm. 7.5.
@@
(a) Let $G$ denote the input quasigroup (of order $n$). First, observe that every quasigroup has a generating set of size $\leq\lceil\log n\rceil$ [Mil78]. Therefore, we start by guessing a subset $X\subseteq G$ of size at most $\leq\lceil\log n\rceil$ (resp. the size bound given in the input). For the decision version, we use $O(\log^{2}n)$ existentially quantified non-deterministic bits ($\exists^{\log^{2}n}$) to guess a generating sequence. To find a minimum-sized generating sequence, we can enumerate all possible generating sequences in $\mathsf{quasiAC^{0}}$. In the next step, we verify whether $X$ actually generates $G$. This is done by checking for all $g\in G$ (universally verifying $O(\log n)$ bits, $\forall^{\log n}$) whether $g\in\langle X\rangle$, which can be done in $\mathsf{NTIME}(\mathrm{polylog}(n))\subseteq\mathsf{quasiAC^{0}}$ by Theorem 7.7. This concludes the proof of the bound (a) for the decision variant.
(c) If we want to compute the minimum generating set, we have to check whether it is actually of smallest possible size. To do this, in a last step, we use the same technique as above to check that all $Y\subseteq G$ with $\left|\mathinner{Y}\right|<\left|\mathinner{X}\right|$ do not generate $G$.
(b) Every quasigroup has a generating set of size $\leq\lceil\log n\rceil$. Existentially guess the generating set using $\log^{2}n$ bits, then use Lem. 7.11 to compute $\langle X\rangle$ in $\mathsf{SAC}^{1}$. With a log-depth tree of bounded fan-in AND gates, we then check that $\langle X\rangle$ is indeed the whole quasigroup. ∎
@@
We may similarly reduce Quasigroup Isomorphism to Membership for quasigroups. This formalizes the intuition that membership testing is an essential subroutine for isomorphism testing and MGS. In particular, in the setting of $\mathsf{quasiAC^{0}}$, we have that isomorphism testing and MGS reduce to Membership. This latter consequence might seem surprising, as in the setting of groups, Membership belongs to $\mathsf{L}$, while MGS belongs to $\textsf{AC}^{1}(\textsf{L})$ (Thm. 7.1) and it is a longstanding open problem whether Group Isomorphism is in P.
@@ 7.3 MGS for Magmas
The decision variant of Minimum Generating Set for magmas is $\mathsf{NP}$-complete.
This $\mathsf{NP}$-completeness result helps explain the use of Integer Linear Programming in practical heuristic algorithms for the search version of this problem, e.g., [JMV23].
@@
It is straightforward to see that the problem is in $\mathsf{NP}$ by simply guessing a suitable generating set. To show $\mathsf{NP}$-hardness we reduce 3SAT to Minimum Generating Set for magmas. Let $F=\bigwedge_{j=1}^{m}C_{j}$ with variables $X_{1},\dots,X_{n}$ be an instance of 3SAT. Our magma $M$ consists of the following elements:
@@
$C_{j}X$ $=T_{j,j}\text{ if the literal $X$ appears in $C_{j}$}$
$TX_{i}$ $=\overline{X}_{i},$
$T\overline{X}_{i}$ $=X_{i},$
$T_{j,k}T_{k+1,\ell}$ $=T_{j,\ell}.$
all other products are defined as 00.
@@
When $F$ is satisfiable, we claim that $M$ can be generated by precisely $n+m$ elements. Namely, include all $C_{j}$ in the generating set. Fix a satisfying assignment $\varphi$ to $F$. If $\varphi(X_{i})=1$, then include $X_{i}$ in the generating set, and if $\varphi(X_{i})=0$, include $\overline{X}_{i}$ in the generating set. Since $\varphi$ is a satisfying assignment, for each $j$, there is a literal $X$ in our generating set that appears in $C_{j}$, and from those two we can generate $C_{j}X=T_{j,j}$. Next, $T_{j,j}T_{j+1,j+1}=T_{j,j+1}$, and by induction we can generate all $T_{j,k}$. In particular, we can generate $T=T_{1,m}$, and then using $T$ and our literal generators, we can generate the remaining literals.
Conversely, suppose $M$ is generated by $n+m$ elements; we will show that $F$ is satisfiable. As argued above, those elements must consist of $\{C_{j}:j\in[m]\}$ together with precisely one literal corresponding to each variable. Since the final defining relation can only produce elements $T_{j,\ell}$ with $j$ strictly less than $\ell$, the only way to generate $T_{j,j}$ from our generating set is using the first relation. Namely, at least one of the literals in our generating set must appear in $C_{j}$. But then, reversing the construction of the previous paragraph, the literals in our generating set give a satisfying assignment to $F$. ∎

### v4

@@ 7.2 MGS for Quasigroups
    The decision version belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{DSPACE}(\log^{2}n)$;
@@
    The search version belongs to $\mathsf{quasiAC^{0}}\cap\mathsf{DSPACE}(\log^{2}n)$.
@@
    not $\exists^{\log^{2}n}\mathsf{P}$-complete under $\mathsf{quasiAC^{0}}[p]$ Turing reductions, or even such reductions up to depth $o(\log n/\log\log n)$.
@@
(a) Thm. 7.5 (a) for MGS, resp. Thm. 6.1 for Quasigroup Isomorphism, place both problems into classes that are contained in $\mathsf{quasiAC^{0}}$ by Fact 2.4. If either problem were $\exists^{\log^{2}n}\mathsf{P}$-complete under quasi-polynomial size, $o(\log n/\log\log n)$-depth, $\mathsf{AC}[p]$ circuits, then, since $\textsc{Parity}\in\mathsf{P}\subseteq\exists^{\log^{2}n}\mathsf{P}$, we would get such circuits for Parity, which do not exist [Raz93, Smo87] (Smolensky’s argument yields a minimum size of $\exp(\Omega(n^{1/(2d)}))$ for depth-$d$ circuits, which is super-polynomial when $d\in o(\log n/\log\log n)$).
(b) Both MGS for quasigroups and Quasigroup Isomorphism are in $\mathsf{DSPACE}(\log^{2}n)$ by Thm. 7.5, resp. [CTW13]. The closure of $\mathsf{DSPACE}(\log^{2}n)$ under poly-log space reductions is contained in $\mathsf{polyL}=\bigcup_{k\geq 0}\mathsf{DSPACE}(\log^{k}n)$. If either of these two quasigroup problems were complete for $\exists^{\log^{2}n}\mathsf{P}$ under polylog-space Turing reductions, we would get $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{polyL}$. Under the latter assumption, by a straightforward padding argument, we now show that $\mathsf{EXP}=\mathsf{PSPACE}$.
@@
Now we return to establishing the main result of this section, Thm. 7.5. To establish Thm. 7.5 (a) and (b), we will crucially leverage the Membership for quasigroups problem. To this end, we will first establish the following.
@@
Membership for quasigroups belongs to $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
@@
The reachability lemma and proofs.
The proofs of Theorem 7.7 and Theorem 7.5 rely crucially on the following adaption of the Babai–Szemerédi Reachability Lemma [BS84, Theorem 3.1] to quasigroups. We first generalize the notion of a straight-line program for groups [BS84] to SLPs for quasigroups. We follow the same strategy as in the proof of [BS84, Theorem 3.1], but there are some subtle, yet crucial, modifications due to the fact that quasigroups are non-associative and need not posses an identity element.
Let $X$ be a set of generators for a quasigroup $G$. We call a sequence of elements $g_{1},\ldots,g_{\ell}\in G$ a straight-line program (SLP for short) if each $g_{i}$ ($i\in[\ell]$) either belongs to $X$, or is of the form or $g_{j}g_{k}$, $g_{j}\backslash g_{k}$, or $g_{j}/g_{k}$ for some $j,k<i$ (where $g_{j}\backslash g_{k}$, resp. $g_{j}/g_{k}$, denotes the quasigroup division as defined in Section 2.1). An SLP is said to compute or generate a set $S$ (or an element $g$) if $S\subseteq\{g_{1},\dots,g_{\ell}\}$ (resp. $g\in\{g_{1},\dots,g_{\ell}\}$).
For any sequence of elements $z_{1},\dotsc,z_{k}$, let $P(z_{1}z_{2}\dotsb z_{k})$ denote the left-to-right parenthesization, e.g., $P(z_{1}z_{2}z_{3})=(z_{1}z_{2})z_{3}$. For some initial segment $z_{0},z_{1},\ldots,z_{i}$ define the cube
$K(i)=\{P(z_{0}z_{1}^{e_{1}}\cdots z_{i}^{e_{i}}):e_{1},\ldots,e_{i}\in\{0,1\}\},$
where $e_{j}=0$ denotes omitting $z_{j}$ from the product (since there need not be an identity element). Define $L(i)=K(i)\backslash K(i)=\{g\backslash h:g,h\in K(i)\}$.
Note that, if $L(k)=G$, then $z_{1},\dotsc,z_{k}$ is very similar to a cube generating sequence and we call it a cube-like generating sequence (the difference is that $L(k)$ allows quotients of cube words, not only cube words, and in a quasigroup quotients of cube words need not always be cube words).
@@
Let $G$ be a finite quasigroup and let $X$ be a set of generators for $G$. Then there exists a sequence $z_{0},\dots,z_{t}$ with $t\leq\log\left|\mathinner{G}\right|$ such that:
  1. (1)
    $L(t)=G$
  2. (2)
    for each $i$ we have either $z_{i}\in X$ or $z_{i}\in\{gh,\,g/h,\,h\backslash g\mid g,h\in L(i-1)\}$.
In particular, for each $g\in G$, there exists a straight-line program over $X$ generating $g$ which has length $O(\log^{2}|G|)$.
@@
  • Otherwise, $L(i)$ is not a sub-quasigroup; hence, there exist $g,h\in L(i)$ with one of $gh$, $g/h$, or $g\backslash h\not\in L(i)$. We choose $z_{i+1}$ to be one of these $gh$, $g/h$, or $g\backslash h\not\in L(i)$.
Next, we first claim that $|K(i+1)|=2\cdot|K(i)|$. Note that $K(i+1)=K(i)\cup K(i)z_{i+1}$ by definition. As right-multiplication by a fixed element is a bijection in a quasi-group, it suffices to show that $K(i)\cap K(i)z_{i+1}=\emptyset$. So, suppose that there exists some $a\in K(i)\cap K(i)z_{i+1}$. Then $a=gz_{i+1}$ for some $g\in K(i)$. Hence, $z_{i+1}=g\backslash a$, contradicting $z_{i+1}\not\in L(i)$, since both $a$ and $g$ are in $K(i)$.
It now follows that $\left|\mathinner{K(i)}\right|=2^{i}$ and, hence, for some $t\leq\lceil\log_{2}(|G|)\rceil$ we have $L(t)=G$. This completes the proof of the first part of the lemma.
[…]
Remark 7.10.
The proof of Lemma 7.9 shows that for any quasigroup $Q$ and any generating set $S\subseteq Q$, every element $g\in Q$ can be realized with a parse tree of depth $O(\log^{2}|Q|)$. In contrast, Wolf [Wol94] establishes the existence of parse trees of depth $O(\log|Q|)$.
For proving Theorem 7.7 we follow essentially the ideas of [Fle22] (though we avoid introducing the notion of Cayley circuits); however, we have to use a more refined approach by not simply guessing an SLP but the complete information given in Lemma 7.9. Fleischer obtained a $\mathsf{quasiAC^{0}}$ bound for Membership for groups by then showing that the Cayley circuits for this problem can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit. We will instead directly analyze the straight-line programs using an $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ (sequential) algorithm.
@@
To decide whether $g\in\langle X\rangle$ with $X=\{x_{1},\dots,x_{\ell}\}$, we guess a sequence of elements $z_{1},\dots,z_{k}\in G$ with $k\leq\log n$ as in Lemma 7.9 together with the necessary information to verify condition (2) in the lemma and whether $g\in L(k)$. This information consists of sequences $\ell_{i}\in[\ell]$ and $e_{i,j}^{(\mu)}\in\{0,1\}$ for $i\in[k]$, $j\in[i-1]$, $\mu\in[4]$ and, as witness that $g\in L(k)$, a sequence $f_{j}^{(\mu)}\in\{0,1\}$ for $j\in[k]$, $\mu\in[2]$. Note that all this information can be represented using $O(\log^{2}n)$ bits.
In the $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ computation we compute for all $i$ the elements $g_{i}^{(\mu)}=P(z_{0}z_{1}^{e_{i,1}^{(\mu)}}\cdots z_{i-1}^{e_{i,i-1}^{(\mu)}})$ and verify that $z_{i}$ is either $x_{\ell_{i}}$ or can be written as $ab$, $a\backslash b$ or $a/b$ where $a=g_{i}^{(1)}\backslash g_{i}^{(2)}$ and $b=g_{i}^{(3)}\backslash g_{i}^{(4)}$ – thus, verifying the conditions of Lemma 7.9. Finally, with the same technique we verify that $g\in L(k)$ by checking whether $g=P(z_{0}z_{1}^{f_{1}^{(1)}}\cdots z_{k}^{f_{k}^{(1)}})\backslash P(z_{0}z_{1}^{f_{1}^{(2)}}\cdots z_{k}^{f_{k}^{(2)}})$. ∎
@@
(a) Let $G$ denote the input quasigroup (of order $n$). First, observe that every quasigroup has a generating set of size $\leq\lceil\log n\rceil$ [Mil78]. Therefore, we start by guessing a subset $X\subseteq G$ of size at most $\leq\lceil\log n\rceil$ (resp. the size bound given in the input). For this, we use $O(\log^{2}n)$ existentially quantified non-deterministic bits ($\exists^{\log^{2}n}$). Furthermore, we also guess some additional information, namely, the sequence $z_{1},\dots,z_{k}\in G$ from Lemma 7.9 and, just as in the proof of Theorem 7.7, the witnesses that this is, indeed, a cube-like sequence. As in the proof of Theorem 7.7 this takes again $O(\log^{2}n)$ existentially quantified bits.
In the next step, we verify whether $X$ actually generates $G$. This is done by checking for all $g\in G$ (universally verifying $O(\log n)$ bits, $\forall^{\log n}$) whether $g\in\langle X\rangle$, which can be done in $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$ just as in Theorem 7.7. Here it is crucial to note that, as we have already guessed the cube-like generating set in the outer existential block, it suffices to guess $O(\log n)$ bits for the exponents. Finally, as in the proof of Theorem 7.7, we verify in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ using the additional information guessed above that $z_{1},\dots,z_{k}\in G$ is indeed, a cube-like sequence. This concludes the proof of the bound (a) for the decision variant.
(b) To find a minimum-sized generating sequence, we can enumerate all possible generating sets $X$ of size at most $\log n$ in $\mathsf{quasiAC^{0}}\cap\mathsf{DSPACE}(\log^{2}n)$. If we want to compute the minimum generating set, we first, have to find a generating set $X$ and then we have to check whether $X$ is actually of smallest possible size. To do this, in a last step, we use the decision variant from above to check that there is no generating set of size at most $\left|\mathinner{X}\right|-1$ for $G$. ∎
@@
While it is possible to directly reduce MGS to Membership in $\mathsf{quasiAC^{0}}$, we obtain slightly better bounds by instead directly leveraging the Reachability Lemma 7.9. It is also possible to $\mathsf{quasiAC^{0}}$-reduce Quasigroup Isomorphism to Membership. This formalizes the intuition that, from the perspective of $\mathsf{quasiAC^{0}}$, membership testing is an essential subroutine for isomorphism testing and MGS. This might seem surprising, as in the setting of groups, Membership belongs to $\mathsf{L}$, while MGS belongs to $\textsf{AC}^{1}(\textsf{L})$ (Thm. 7.1), yet it is a longstanding open problem whether Group Isomorphism is in P.
@@ 7.3 MGS for Magmas
The decision variant of Minimum Generating Set is $\mathsf{NP}$-complete under many-one $\mathsf{AC^{0}}$ reductions for commutative magmas (unital or not).
This $\mathsf{NP}$-completeness result helps explain the use of Integer Linear Programming in practical heuristic algorithms for the search version of this problem in magmas, e. g., [JMV23].
@@
It is straightforward to see that the problem is in $\mathsf{NP}$ by simply guessing a suitable generating set. To show $\mathsf{NP}$-hardness we reduce 3SAT to Minimum Generating Set for commutative magmas. Let $F=\bigwedge_{j=1}^{m}C_{j}$ with variables $X_{1},\dots,X_{n}$ and clauses $C_{1},\dots,C_{m}$ be an instance of 3SAT. Our magma $M$ consists of the following elements:
@@
$C_{j}X$ $=S_{j,j}\text{ if the literal $X$ appears in $C_{j}$}$
$SX_{i}$ $=\overline{X}_{i},$
$S\overline{X}_{i}$ $=X_{i},$
$S_{j,k}S_{k+1,\ell}$ $=S_{j,\ell}.$
Aside from multiplication being commutative (e. g., we also have $X_{i}T=\overline{X}_{i}$, etc.), all other products are defined as 00.
The idea is that the presence of $S_{j,k}$ in a word indicates that clauses $j,j+1,\dotsc,k$ have been satisfied. This interpretation aligns with the multiplication above: viz. $C_{j}X=S_{j,j}$ if $X$ satisfies $C_{j}$. If clauses $j$ through $k$ are satisfied ($S_{j,k}$) and clauses $k+1$ through $\ell$ are satisfied ($S_{k+1,\ell}$), then in fact clauses $j$ through $\ell$ are satisfied ($S_{j,k}S_{k+1,\ell}=S_{j,\ell}$). We use “$S$” for all of these as a mnemonic for “satisfied” and also because $S=S_{1,m}$ acts as a “swap” on literals.
@@
When $F$ is satisfiable, we claim that $M$ can be generated by precisely $n+m$ elements. Namely, include all $C_{j}$ in the generating set. Fix a satisfying assignment $\varphi$ to $F$. If $\varphi(X_{i})=1$, then include $X_{i}$ in the generating set, and if $\varphi(X_{i})=0$, include $\overline{X}_{i}$ in the generating set. Since $\varphi$ is a satisfying assignment, for each $j$, there is a literal $X$ in our generating set that appears in $C_{j}$, and from those two we can generate $C_{j}X=S_{j,j}$. Next, $S_{j,j}S_{j+1,j+1}=S_{j,j+1}$, and by induction we can generate all $S_{j,k}$. In particular, we can generate $S=S_{1,m}$, and then using $S$ and our literal generators, we can generate the remaining literals.
Conversely, suppose $M$ is generated by $n+m$ elements; we will show that $F$ is satisfiable. As argued above, those elements must consist of $\{C_{j}:j\in[m]\}$ together with precisely one literal corresponding to each variable. Since the final defining relation can only produce elements $S_{j,\ell}$ with $j$ strictly less than $\ell$, the only way to generate $S_{j,j}$ from our generating set is using the first relation. Namely, at least one of the literals in our generating set must appear in $C_{j}$. But then, reversing the construction of the previous paragraph, the literals in our generating set give a satisfying assignment to $F$.
Identity elements. We may add a new element $e$, relations $ea=ae=a$ for all $a$, and include $e$ as a generator. When this is done, the reduction queries whether the algebra is generated by $n+m+1$ elements or more than that many, since the element $e$ must be contained in every generating set. ∎